### v1

# Extraction of Medication and Temporal Relation from Clinical Text by Harnessing Different Deep Learning Models

Abstract
Clinical texts, represented in electronic medical records (EMRs), contain rich medical information and are essential for disease prediction, personalised information recommendation, clinical decision support, and medication pattern mining and measurement. Relation extractions between medication mentions and temporal information can further help clinicians better understand the patients’ treatment history. To evaluate the performances of deep learning (DL) and large language models (LLMs) in medication extraction and temporal relations classification, we carry out an empirical investigation of MedTem project using several advanced learning structures including BiLSTM-CRF and CNN-BiLSTM for a clinical domain named entity recognition (NER), and BERT-CNN for temporal relation extraction (RE), in addition to the exploration of different word embedding techniques. Furthermore, we also designed a set of post-processing roles to generate structured output on medications and the temporal relation. Our experiments show that CNN-BiLSTM slightly wins the BiLSTM-CRF model on the i2b2-2009 clinical NER task yielding 75.67, 77.83, and 78.17 for precision, recall, and F1 scores using Macro Average. BERT-CNN model also produced reasonable evaluation scores 64.48, 67.17, and 65.03 for P/R/F1 using Macro Avg on the temporal relation extraction test set from i2b2-2012 challenges. Code and Tools from MedTem will be hosted at https://github.com/HECTA-UoM/MedTem
1 Introduction
With the continuous growth of the need for medical text mining, extraction of temporal information has become a significant field of Natural Language Process (NLP) Su et al. (2004), e.g. when the disease was diagnosed, how long a patient should take some medicines. Due to its importance for many other NLP tasks, temporal information tagging and annotation from clinical texts have become an active research field (Figure 1 for example). Clinical texts including electronic health records (EHRs) consist of medical history, diagnoses, medications, treatment plans, immunisation dates, allergies, radiology images, and laboratory and test results, which contain a lot of useful information that can provide the opportunity to support medical research Han et al. (2023). These include the study of the side effects of medications (a.k.a adverse drug effects) taken over time, helping doctors make better decisions for patients, and finding the correlation between drug use and its outcomes, etc. Fredriksen et al. (2014); Alrdahi et al. (2023).
Figure 1: Original Text Example from EMRs
In view of the importance of temporal information in clinical free texts, in order to extract it from unstructured data and convert it to structured features consisting of temporal relations (TLINKs) between events and time expressions, researchers have developed some techniques, tools, and workflows. Generally, they can be divided into three main categories: rule-based approaches, machine learning and combined approaches. Most machine learning methods rely on feature-engineered supervised learning to predict temporal relations between events, which require tagged datasets that cost a lot of effort and time of experts in the medical domain.
With the development of various deep learning (DL) models nowadays, this work aims at an empirical investigation on how they perform on this specific challenge, i.e. medication and temporal relation extraction from clinical text.
There are two major sub-tasks: 1) Medication Entity Recognition, which involves finding and identifying named drugs mentioned in unstructured text, the first step toward automating the extraction of medications from EMRs. 2) Automatic Temporal Relation Extraction from free text EMRs. We approach these two steps by applying several advanced DL methods including the BiLSTM+CRF, CNN-BiLSTM, and BERT-base-CNN models to train on the available data sets. Additionally, a SparkNLP tool (DateNormalizer) Kocaman and Talby (2021) is used to extract and normalise dates from relative date-time phrases.
2 Methodological Design
Our designed methodology includes five main steps as shown in Figure 2, which are Data Pre-processing, Word Embedding, Modelling, Time Expression and Event Extraction, and Temporal Relations and Medications Candidates Classification. The Step-1 on data pre-processing includes sentence segmentation, tokenisation, Part of Speech (POS) tagging, and syntactic parsing, which details are located in Appendix with examples (A: Pre-Processing Details: “A: Pre-Processing Details”). The goal of data pre-processing is to produce structured text with features. Regarding Step-2, we tried three different kinds of word embedding techniques including one-hot encoding, bag-of-words, and GloVe embedding, which are detailed in Appendix (B: Word Embeddings: “B: Word Embeddings”).
Figure 2: Methodology Illustration on Overall Tasks of MedTem
2.1 Deployed Learning Models
2.1.1 BiLSTM
In machine learning fields, the problem that the model cannot get the previous moment information due to the vanishing gradient is also called the long-term dependence problem. In order to solve this problem, many researchers have tried to improve the Recurrent Neural Network (RNN) model, including BRNN Arisoy et al. (2015), GRU Tang et al. (2016), and LSTM Hochreiter and Schmidhuber (1997), etc., among which the LSTM model is the most widely used one.
LSTM is an improved RNN structure that can effectively avoid the long-term dependence problem that the traditional RNNs cannot solve. In the later work, many people adjusted and popularised the model, and the modern, widely used LSTM structure came into being. LSTM network has the same chain structure as traditional RNNs, but the repeating module has a different design. The RNN model only has a single layer of repeated modules. In contrast, the LSTM network has four interacting neural network layers, which makes the LSTM model more complex and sacrifices a certain amount of computational time, but the model output becomes better. The standard LSTM model is shown in Figure 7 (Appendix C: More Details on Learning Models), which mainly includes three groups of adaptive element multiplication gates, namely, the input gate, forgetting gate, and output gate.
In actual text processing applications, the text is characterised by sequential feature correlation, and words in the current moment are not only affected by words in the past moment but also by words in the future moment. In order to make the model contain contextual information at the same time, researchers put forward a bidirectional LSTM model (BiLSTM) Zhang et al. (2015); Zhou et al. (2016). BiLSTM contains two LSTM layers, which train the forward and backward sequences, respectively, and both LSTM layers are connected to the output layer. This forward-backward two-layer LSTM structure can provide historical information and future information to the output layer at the same time, which can not only retain the advantages of LSTM to solve long-term dependence but also take into account the context information and effectively deal with sequence problems.
2.1.2 CRF model
Conditional random fields (CRFs) Lafferty et al. (2001) are a type of statistical modelling methodology frequently employed in pattern recognition and machine learning, as well as for structured prediction, which is a widely-applied modelling strategy for many NLP tasks Han et al. (2013, 2015). Data having an underlying graph structure can be better modelled using CRFs, a graphical model that can take advantage of structural dependencies between outputs. In contrast to classifiers, which make label predictions for isolated data without considering their “neighbouring” samples, CRFs are able to account for context Maldonado et al. (2017); Moreau et al. (2018); Wu et al. (2022). In order to accomplish this goal, the predictions are modelled in the form of a graphical model, which depicts the existence of relations between the predictions. The choice of the graph to implement is context-specific. In NLP fields, for instance, “linear chain” CRFs are prevalent because each prediction depends only on its immediate neighbours. Connecting neighbouring and/or analogous regions ensures that all regions within a picture are predicted in the same way by the graph. The mathematical calculation of CRFs is detailed in Appendix (C: More Details on Learning Models).
2.1.3 CNN model
In deep learning (DL), a classic convolutional neural network (CNN) is a class of artificial neural networks (ANN), most commonly applied to analyse visual imagery Albawi et al. (2017). A CNN model mainly consists of these layers: input layer, convolution layer, relu layer, pooling layer and full connection layer. One of CNN’s most alluring quality is its ability to make use of spatial or temporal correlation in data. CNN is divided into multiple learning phases, each of which is made up of a combination of convolutional layers, nonlinear processing units, and subsampling layers Goodfellow et al. (2016). In a CNN, each layer of the network uses a set of convolutional kernels to perform many transformations. The convolution process simplifies extracting useful features from spatially correlated data points. More on CNNs are displayed in Appendix C.
2.1.4 BERT model
Brief for “Bidirectional Encoder Representations from Transformers”, BERT Devlin et al. (2018) is founded on Transformer structure Vaswani et al. (2017), a cutting-edge DL model in which all output elements are related to all input elements and the weights between them are dynamically computed based on their connections. BERT is unique in that it can be used with text that is read either from left to right or right to left. With this characteristic, BERT was originally taught for two NLP tasks: Next Sentence Prediction and Masked Language Modelling. The input for BERT is comprised of Token Embeddings, Segment Embeddings, and Position Embeddings, followed by pre-training as indicated in the figure 11 (Appendix).
The final output of the [CLS] parameter of the BERT is set as C, the input embedding is designated as E, and the final output of the ith token is designated as T. This will enable effective dynamic feature encoding of polysemous words. In other words, the same word might produce a variety of word vector outputs depending on the language environment. Additionally, the pre-trained model’s embedding layer transfers the knowledge from the corpus to make the word vectors more generic, greatly enhancing the model’s accuracy.
Instead of using word embedding, the conventional method of language processing, BERT assigns each word to a vector that only conveys a single dimension of its meaning. Extensive quantities of labeled data are required for word embedding models. However, because all words are in some sense bound to a vector or meaning, they have difficulty with the context-heavy, predictive nature of question responding. In order to prevent the target word from “seeing itself” (i.e., developing a meaning apart from its context), BERT employs a technique called masked language modelling. The advent of pre-trained language models in recent years has marked a significant milestone in developing NLP. Pre-trained language model methods from the past focused on acquiring high-quality word embeddings. Word vectors are context-free, notwithstanding their ability to capture a word’s semantics. Since this is the case, individuals are unable to understand the deeper meanings included in the text, such as its syntax and semantics.
2.2 Time Expressions and Events Named Entity Recognition
The goal of NER task is to mark up temporal information present in clinical text in order to enable reasoning on relevant events (medications) for each patient. In this work, a commonly used format in entity tagging “Inside–outside–beginning (IOB)” Ramshaw and Marcus (1999), is used, which is explained below:
  1. 1.
    the B-prefix indicates that the tag is at the beginning of a chunk that follows another chunk without O tags between the two chunks.
  2. 2.
    the I-prefix indicates that the tag is inside a chunk.
  3. 3.
    the O-prefix indicates that the token belongs to no chunk.
In this step, a BiLSTM+CRF model and a CNN-BiLSTM model are deployed to achieve the goal and a comparison of performance is taken to demonstrate the difference, advantages, and disadvantages between them.
2.3 Temporal Relations and Medication Candidates Classification
Relation extraction is the task of predicting attributes and relations for entities in a sentence. Extracted relations usually occur between two or more entities of a particular type (e.g. Medication, Dosage) and fall into several semantic categories. In this work, these relations that are discussed are between medications and corresponding dates (e.g. “He has been treated with simvastatin since 12/04/1998”, which contains the relation “simvastatin” happened after “12/04/1998”). Generally, there are eight types of temporal relations, including before, after, simultaneous, overlap, begun_by, ended_by, during, and before_overlap suggested by Sun et al. (2012). However, only the types of before, after, and overlap meet our task. A BERT-based model with CNN as an additional learning layer is implemented for this step.
2.4 Post-processing
After the phrases of entity recognition and relation extraction, to make the result more robust and representative, this post-processing step is to convert results to a structured table or CSV shape. However, in the previous 2012 i2b2 NLP challenge on temporal relation extraction, post-processing was considered a set of classification routines for the EVENTs only, and the TIMEx catergory remained unclassified Lin et al. (2013). In this work, we apply an open resource tool SparkNLP [^1] https://github.com/JohnSnowLabs/spark-nlp, combined with our medication entity recognition and temporal relation extraction system to generate more informative results and a structured representation table e.g. the example Table 1 generated from the example text in Figure 1.
ID Event Statues Start Stop
134529565 Methotrexate ON May 2019 February 2020
134529566 Methotrexate OFF February 2020 Unknown
Table 1: Extraction Result Example
3 Experimental Evaluations
3.1 Dataset and Evaluation Metrics
In this work, the i2b2 challenge track 2009 and 2012 datasets were used. The objective is to gather details about every drug that the patient is known to take for each patient report that is given. Partners Healthcare’s discharge summaries will be the input for the medication challenge Sun et al. (2012). These files are loosely organised. Many drugs are covered in the narrative text in addition to being listed by name in lists of pharmaceuticals. According to statistical analysis, it contains 447 EMRs, with 252 annotated, which could be used to train and test the model. While the data from the i2b2-2012 temporal relations challenge has similar content as the challenge track of 2009 does, it has a different format, additional entity categories, and temporal relations between events. It consists of 760 files including 190 TXT files (texts), 190 XML.EXTENT files (annotation and position in original text), XML files (annotation and text) and TLINK files (relations) Gurulingappa et al. (2012). The data sets are divided into 70%, 15% and 15% for training, validation, and testing, respectively on both these two i2b2 challenges. The evaluation metrics we used are Precision, Recall, F1 score, and Accuracy based on exact matching, detailed in Appendix (“E: Dataset and Metrics”) together with dataset analysis.
3.2 Model-I: BiLSTM-CRF for NER Task
Combined with the two models BiLSTM and CRF, we carried out the NER task of electronic diseases using BiLSTM-CRF. The specific steps carried out are as follows:
  1. 1.
    Apply the Glove or bag-of-word model to obtain words’ semantic representation.
  2. 2.
    Conduct sequence learning using the BiLSTM network, which effectively avoids the gradient disappearance and explosion of other DL models.
  3. 3.
    Restrict the sequence relationship between tags by using CRFs.
Using one-hot word embedding, the evaluation scores on 2009 and 2012 data sets are shown in Table 2 and 3 with the explanations of each tag in Table 9 and 10. The parameter sets for these two experiments as well as the learning curves on Accuracy and Loss changes are listed in Appendix (F: Model Analysis).
Table 2: NER Evaluation on the 2009 Dataset with BiLSTM-CRF.
category Number / % Number
precision recall f1-score support
PADDING 99.99 1.00 1.00 111422
B-do 81.88 70.52 75.78 519
I-do 84.09 74.66 79.10 446
B-m 81.35 66.51 73.19 1069
I-m 87.60 70.42 78.07 622
B-f 83.50 77.48 80.37 444
I-f 75.44 61.43 51.81 166
B-du 61.11 15.28 24.44 72
I-du 44.44 17.78 25.40 135
B-r 39.80 20.42 26.99 191
I-r 41.67 10.64 16.95 141
B-mo 83.05 78.61 80.77 374
I-mo 0.00 0.00 0.00 20
o 95.87 98.50 97.17 32429
accuracy 98.63 148050
macro avg 68.56 53.76 58.55 148050
weighted avg 98.46 98.63 98.50 148050
Table 3: NER Evaluation on the 2012 Dataset with BiLSTM-CRF.
category Number / % Number
precision recall f1-score support
PADDING 1.00 99.99 99.99 63867
B-CLINICAL_DEPT 68.85 32.81 0.44 128
I-CLINICAL_DEPT 80.43 54.95 65.29 202
B-EVIDENTIAL 1.00 4.11 7.89 73
I-EVIDENTIAL 0.00 0.00 0.00 6
B-OCCURRENCE 61.69 44.39 51.63 410
I-OCCURRENCE 34.69 7.46 12.27 228
B-PROBLEM 59.10 56.64 57.85 602
I-PROBLEM 75.67 61.92 68.11 864
B-TREATMENT 66.80 68.31 67.55 486
I-TREATMENT 65.87 67.63 66.74 448
B-TEST 58.48 54.18 56.25 299
I-TEST 64.42 63.23 63.82 378
O 87.82 95.83 91.65 75.82
accuracy 97.11 75573
macro avg 65.99 50.82 53.82 75573
weighted avg 96.89 97.11 96.88 75573
As mentioned, evaluation metrics were based on exact-matching criteria and with this setting, the BiLSTM+CRF model achieves a 0.98 weighted average accuracy and 0.69 of Macro averaging. Apparently, combined with the evaluation scores of each subcategory, it can be seen that the amount of entities of each category varies significantly due to the uneven data distribution, which leads to the scores between entity categories varying wildly.
As demonstrated in the result of 2009 data, this BiLSTM+CRF model has better prediction performance on class “o” and “B-m” in the 2009 dataset. In general, the BiLSTM+CRF model shows better performance on the 2009 dataset than the 2012 dataset for the NER task.
In order to explore the impacts of different embedding inputs for the BiLSTM+CRF model, we trained the model with Glove (glove.6B.50d), word2vector, in addition to one-hot encoding, and the average and standard deviations of metrics are recorded. Table 4 demonstrates the performance of each model on the same dataset. Glove embedding achieved 72.48% of precision, 52.58% of recall, and 60.86% of f1-score, respectively which increased precision by around 2%. For word2vec embedding, this model achieved exact 73.62% precision, 60.34% of recall, and 61.48% of f1-score achieving an further improvement of 1.14% in precision, an increase of 7.76% in recall and an improvement of 0.62% than Glove embedding.
Table 4: Macro Avg Results of Different Embeddings on the 2009 Test Set using BiLSTM+CRF.
embedding precision recall f1-score
one-hot 68.56 53.76 58.55
Glove 72.48 52.58 60.86
word2vec 73.62 60.34 61.48
Table 5: Macro Avg Results of Experiments on the 2009 Test Dataset with CNN-BiLSTM.
model precision recall f1-score
CNN-BiLSTM 75.67 77.83 78.17
For the overall trend, we observed that the prediction performance really depends on the amount of sub-category data for training, which is that the more data, the more information obtained, leading to a better evaluation score. For entity categories with a smaller number of entries, the recognition effect is poorer due to insufficient model learning. For example, sub-classes B-r and I-du have lower precisions than B-m and B-do in the 2009 dataset. However, there are some exceptions that with more training data samples, the model still shows poorer results for sub-class B-r than sub-class B-du.
In addition, the results based on precision, recall, and F1-score are pretty different. For most sub-classes, the model has higher precision than recall values which means it returns fewer results, but most of its predicted labels are correct when compared to the gold labels. Namely, the percentage of Negative labels is lower. In conclusion, the BiLSTM+CRF model can not fully solve the problem of unbalanced data distribution. On the dataset with unbalanced data distribution, the performance for the entity category with a smaller number of samples is lower.
3.3 Model-II: CNN-BiLSTM for NER
The Convolutional Neural Network (CNN) used in this approach draws inspiration from Chiu and Nichols (2016) hybrid model, which uses bidirectional LSTMs to learn both character and word-level information. As a result, our model employs word embeddings, in addition to other human-created word features and character-level information retrieved using a convolutional neural network. All of these features are sent into a BiLSTM for each word token. Following the tutorial by Hofer (2018), we adapted the code from Chiu and Nichols (2016), to implement the CNN-BiLSTM model for NER task on the 2009 dataset.
The idea of this model is to perform NER prediction by CONCAT of both character-level and word-level features and their combined input knowledge into BiLSTM learning. This enables the model to make better use of character-level features such as prefixes and suffixes, which can reduce the work of manual feature construction. The character-level features are obtained by a CNN model which has been approved as a great success in the NER task and POS tagging task Chotirat and Meesad (2021); Labeau et al. (2015). Per-character feature vectors, such as character embeddings and character kinds, are concatenated with a max layer to produce a new feature vector, which is then used to train a model for recognising words.
The BiLSTM layer forms the core of the network and has the following three inputs: character-level patterns identified by CNN, word-level input from GloVe embeddings, and casing input (word case changes). One of the critical tasks here is feature fusion as mentioned by Chiu and Nichols (2016). Additional features can be applied for example, words can be divided into six categories, including all uppercase letters, all lowercase letters, all numbers, some numbers, and so on. The model includes 7 layers of network structure: character embedding, dropout, 1D convolution, word embedding, concatenation, BiLSTM, and dense out layer (ref. Appendix F: Model Analysis). This CNN-BiLSTM model was set for training for 30 Epoch but convergence after 14 Epoch. Table 5 presents the performance (macro average precision, recall and f1-score) of this model which is better than the scores from the BiLSTM-CRF model. Table 13 shows the parameters used in this model (Appendix).
3.4 Model-III: BERT-CNN for Temporal Relation Extraction
Although the [CLS] tags in BERT output can achieve reasonable classification results, the rich semantic knowledge contained in BERT is not fully utilised. Therefore, in this work, inspired by Michalopoulos et al. (2020), the CNN model is fused to expand BERT learning (structure demonstrated in Figure 24). After the BERT model receives the processed text, the content of the EMRs text is represented by a vector through the two-layer Transformer mechanism of the BERT model. The model produces a vectorised representation of comprehensive semantic information fused by each token vector, sentence vector, and feature vector in the EMRs and then feeds the output of the model to the convolutional neural network. In this work, three different convolutional kernels are used to capture different feature information. After the fully connected layer is connected by word vector mapping, the CNN model further extracts the semantic information of the dialogue text.
3.4.1 Sampling and Evaluation
In order to mitigate the influence of the training dataset imbalance of each sub-class, we down-sampled the dataset to 3000 samples for each class (AFTER, OVERLAP and BEFORE). Table 6 gives the result of the prediction on the training dataset. Table 7 shows the result of $BERT_{base}+CNN$ model performance for the relation extraction (RE) task on the i2b2-2012 test data. This model achieves a macro avg. 64.48% of precision, 67.17% of recall, and 65.03% of F1-score.
Table 6: Temporal RE Evaluation on the 2012 Sampled Training Dataset using BERT-CNN.
category Number / % Number
precision recall f1-score support
AFTER 96.23 97.74 96.86 3000
OVERLAP 93.71 95.85 94.80 3000
BEFORE 97.13 93.24 95.06 3000
accuracy 95.60 9000
macro avg 95.60 95.60 95.60 9000
weighted avg 95.60 95.60 95.60 9000
Table 7: Temporal RE Evaluation on the 2012 Test Dataset using BERT-CNN
category Number / % Number
precision recall f1-score support
AFTER 34.52 49.00 40.52 1122
OVERLAP 67.34 77.23 71.94 4078
BEFORE 91.62 75.28 82.64 6000
accuracy 73.32 11200
macro avg 64.48 67.17 65.03 11200
weighted avg 77.03 73.34 74.47 11200
As the prediction shows, with BERT-base as a word vector extraction or directly as a classification model, the BERT+CNN model can achieve a relatively good classification result, indicating that the pre-trained model BERT-base can extract the semantic information well of dialogue text. Although the BERT+CNN model has reasonable performance for relation extraction, over-fitting on the training dataset is still inevitable, especially for sub-class AFTER and OVERLAP, which show 61.71% and 26.37% lower than the prediction results on the training dataset, respectively. There is still the correlation between the number of test data samples and prediction precision which is that the more supporting samples, the better the prediction performance, which is similar to our NER findings. Finally, our post-processing rules and example outputs are displayed in Appendix (Post-Processing).
4 Related Work
NER and RE belong to the information extraction (IE) task, which has been developed using different methods, from rule-based ones, to machine learning and deep learning models. The rule-based methods require expertise in designing hand-crafted rules suitable for specific tasks and domains. While these methods worked out in some situations, it is often difficult for other researchers to adopt the existing rules into a new experiment Ji et al. (2019). Remarkable work in this category includes Reeves et al. (2012); Strötgen and Gertz (2013); Hao et al. (2018) on temporal expression extractions, and Roberts et al. (2013); Lin et al. (2013) on medical terminology mining.
Earlier stage ML models on TLINKS detection and clinical events and concepts extraction include Chang et al. (2013); Xu et al. (2013); Wu et al. (2015, 2017) where CNNs and RNNs were explored. The combined model of BiLSTM-CRF was also investigated by some existing work reported in Kim and Lee (2020); Lee et al. (2020); Jiang et al. (2019) and in language-specific experiments such as Chinese EMRs by Ji et al. (2019).
Based on these previous findings, our work further investigated various combinations of these state-of-the-art learning models on clinical domain NER and Temporal modelling. More detailed literature can be found in our Appendix (H: More Background and Literatures).
5 Conclusions and Future Work
We carried out empirical study on medication extraction and temporal relation classification using three hybrid machine learning models BiLSTM-CRF, CNN-BiLSTM, and BERT-CNN. By combining these two tasks and a set of rules we designed, we aim to generate a simple structured output indicating if certain medications are in use during the extracted timeframe. While the experimental work demonstrated promising results, we plan to further improve the model performances from different aspects, which include the integration of domain-specific pertained BERT model such as Med-BERT Rasmy et al. (2021), more text features, data-set augmentation e.g. using synthetic data Belkadi et al. (2023), and system optimisation using different learning structures. We also investigate the prompt-based engineering for MedTem2.0 project Cui et al. (2023).
Limitations
The main sources of our experimental data are the i2b2 Challenge 2009 and 2012, which are considerable corpora leading to a lack of computation resources to do more experiments for fine-tuning the best model settings and parameters. Moreover, the proportion of data (those annotated) on the relationship between medications and their corresponding dates is inadequate, which can not provide sufficient information for MEDICATION-DATE relation extraction. Therefore, the performance of information extraction generated by the rules created for patients’ medication usage status falls short of ideal standards, which requires further improvement.
Ethical Statement
There are no ethical concerns in this work since the data sets we used from n2c2 challenges are anonymised and ethically approved by the shared task organisers.
Appendix
A: Pre-Processing Details
Clinical texts are mixed with different formats. The target of pre-processing is trying to format different data with certain rules and enrich texts with their potential features.
Typically, Pre-processing follows 5 steps: Sentence Segmentation, Tokenisation, Parts of Speech (POS, which explains how a word is used in a sentence), Tagging and Parsing. For our work, GATE Cunningham H (2002), cTAKES Savova et al. (2010) and MedSpaCy (a new clinical text processing toolkit in Python) shows excellent power in building rule-based modules to retrieve extra features. Additionally, the library ’spacy’ is valuable and powerful to link mentions to the associated verb.
Original word Lemmatization POS Tag Explain
The the DET DT det
canning canning NOUN NN compound
process process NOUN NN nsubjpass
for for ADP IN prep
food food NOUN NN pobj
was be AUX VBD auxpass
invented invent VERB VBN ROOT
by by ADP IN agent
Nicolas Nicolas PROPN NNP compound
Appert Appert PROPN NNP pobj
by by ADP IN prep
1999 1999 NUM CD pobj
. . PUNCT . punct
Table 8: POS Tagging Example Marcinkiewicz (1994)
Sentence Boundary Disambiguation
The task of sentence boundary detection is to identify sentences within a text which appear first in an NLP task pipeline, and its performance could decide the following pipeline parts Koehn (2009). Typically, the full stop is detected as the end of a sentence, which is obviously ambiguous and not enough since the full stop is not only used as a sentence boundary delimiter but also juxtaposed with abbreviations (e.g. M.r, A.M, Doc.) which appear quite a lot in clinical texts. Besides, full stop occurs in some special expressions including dates (e.g. 2022.07.22), numbers (e.g. 10.0 mm), and web links (e.g. https://www.google.com/). Sometimes, other delimiters like “!”, “?” could be the end of sentences as well.
In this work, statistical sentence segmentation of Spacy is used and refined with custom rules. For example, given sentence:
The patient ’s bilirubin level at 24 hours of life was 4.6.
Spacy returns three sentences:
  1. 1.
    The patient
  2. 2.
    ’s
  3. 3.
    bilirubin level at 24 hours of life was 4.6
Here we want to make sure that “ ’s ” tokens will never be the start of a sentence. So a rule is added to make sure “ ’s ” never occur at a start of a sentence.
Tokenisation
Having paragraphs segmented into sentences, sentences will be split into tokens, also known as Tokenisation, that is to decompose long texts such as sentences, paragraphs, and articles into word-based data structures for the convenience of subsequent processing and analysis Koehn (2009); Manning and Schutze (1999). Text is some unstructured data; we need first to transform these data into structured data, structured data can be converted into mathematical problems, and Tokenisation is the first step of transformation.
Compared to other script languages (e.g. Chinese), Tokenisation is easier for English corpora since tokens could be separated by space. However, in order to preserve integral information of original texts, endings of contractions ( “’re” in “we’re”, “’t” in “doesn’t”) should receive special treatment (added rules).
POS Tagging
POS Tagging, also called grammatical tagging or word category disambiguation (the POS tags from the Penn Treebank project Marcinkiewicz (1994), which are widely used in practice), refers to the process of marking a word in the corpus as corresponding to a particular part of speech, based on both its definition and context and the output of POS Tagging is a good feature for supervised learning as an input which makes it an important step:
  1. 1.
    POS tagging is important for word sense disambiguation.
  2. 2.
    For example, consider the sentence “They can fish”; One interpretation: They/PRON can/AUX fish/VERB; and another: They/PRON can/VERB fish/NOUN
  3. 3.
    POS tagging is of importance in applications such as machine translation and information retrieval.
Syntactic Parsing
In order to understand sentences, we cannot treat each word in isolation. We need to determine what a word is attached or connected to. Syntactic parsingFrazier (1979) is a technique by which segmented, tokenised, and part-of-speech tagged text is assigned a structure that reveals the relationships between tokens governed by syntax rules, e.g. by grammars. In NLP, rules of context-free grammar (CFG) or probabilistic context-free grammar (PCFG) are used to analyse sentences. Building a parser from scratch is a very complex task in itself. We would pick a grammar like CFG or PCFG, then decide upon which type of parser we want to build. Based on that, we would implement specific algorithms to develop our very own parser. For example, given the sentence “The patient was not started on antibiotics”, the result of Spacy parsing visualisation (Figure 3):
Figure 3: Syntactic Parsing Example using Spacy
B: Word Embeddings
Word embedding techniques play a role in capturing the meaning, semantic relationship, and context of different words while creating word representations Pham and Le (2018). A word embedding is a technique used to generate a dense vector representation of words that include context words about their own. In addition to one-hot embedding that results in sparse matrix embedding, there are enhanced versions of simple bag-of-words models, such as word counts and frequency counters, that essentially represent sparse vectors.
One-hot Encoding
A popular encoding technique is called “one-hot encoding”. In NLP, if a dictionary or sequence has N fields, each field can be represented by an n-dimensional one-hot vector. For example, “The patient was not started on antibiotics” is encoded as in Figure 4 as an example. One-hot encoding can convert classified data into a unified digital format, which is convenient for machine learning algorithms to process and calculate since computers cannot understand non-numeric data. And the fixed dimension vector is convenient for the machine learning algorithm to calculate the linear algebra.
Figure 4: An example of mapping a sentence to a matrix with One-hot Encoding
Data that don’t relate to one another can benefit from one-hot encoding. The arrangement of numbers is treated as a significant characteristic by machine learning algorithms. To put it another way, people will perceive a higher number as superior or more significant than a lower one.
Bag-of-words embedding
A bag-of-words model (BoW) is a way of extracting features from the text for use in modelling, such as with machine learning algorithms, which is a popular way of representing documents Shao et al. (2018). Throughout the realm of information retrieval, the BoW model is used in information retrieval on the assumption that a text may be reduced to a collection of words without regard to its actual order within the document, its grammar, or its syntax. Each word’s presence in the document is autonomous and unrelated to the incidence of other terms. In other words, each word at each place in the document is individually selected without regard to the document’s semantics.
For given two sentences,
  1. 1.
    She is admitted from hospital for a coronary artery bypass surgery.
  2. 2.
    She is a female admitted to outside hospital with chest discomfort.
Based on these two sentences, a dictionary that contains 16 words is built: Dictionary = 1:“She”, 2. “is”, 3. “admitted”, 4. “from”, 5. “hospital”, 6. “for”, 7. “a”, 8. “coronary”, 9. “bypass”, 10. “surgery”, 11. “female”, 12. “to”, 13. “outside”, 14. “with”, 15. “chest”, 16. “discomfort” Then the two sentences can be represented as:
  1. 1.
    $1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0$
  2. 2.
    $1,1,1,0,1,0,1,0,0,0,1,1,1,1,1,1$
Each element in the vector represents the number of times related words in the dictionary appear in the document. However, in the process of constructing document vectors, we can see that we do not express the order in which words appear in the original sentence.
GloVe Embedding
GloVe stands for “Global Vectors” Pennington et al. (2014). Compared to Word2vec, which only keeps local statistic information, the advantage of GloVe is that it captures both global statistics and local statistics of a corpus, and it is an unsupervised learning pre-trained model that represents words in sentences. It has the ability to extract semantic relationships. The main idea of the Glove method is that “semantic relationships between words can be extracted from the co-occurrence matrix” Bullinaria and Levy (2007). Given a document having a sentence of n words, the co-occurrence matrix X will be a n*n matrix (as in Figure 5) where $X_{ij}$ indicates how many times word i has co-occurred with the word j.
Figure 5: Co-occurrence Matrix with a Window Size of 1
Let $P_{ij}$ represent the likelihood of seeing the words we and j together, which is calculated by dividing the number of times we and j occurred together ($X_{ij}$) by the total number of times word we appeared in the corpus ($X_{i}$), where $P_{ij}=X_{ij}/X_{i}$. Suppose there is a function F that takes in word vectors of we j, and k and returns the desired ratio.
$F(w_{i},w_{j},u_{k})=P_{i}k/P_{j}k$
C: More Details on Learning Models
BiLSTM Illustration
It is very similar to the structure of the traditional recurrent neural networks and multi-layer feed-forward neural networks after expansion, so the problem of gradient disappearance will inevitably occur if there are too many layers during training Hochreiter (1998). The schematic diagram of gradient disappearance is shown in Figure 6. The node colour depth in the figure indicates how much input information at the first moment can be retained at the current moment. The darker the colour is, the more information is retained, and the better the model effect is; the lighter the colour is, the less information is retained, and the model is prone to gradient disappearance. From the feed-forward process of the model, as time goes by, the information that can be extracted at subsequent moments gradually decreases. As shown in the figure (6), when analysing the data from time step 7, it is practically impossible to find the information from time step 1 anymore. When the error of the output layer at time step 7 propagates forward through the gradient during the training process, the model is unable to successfully update the relevant weights at the earlier time because of the gradient’s decline Jiang et al. (2019).
Figure 6: The Vanishing Gradient Problem in Deep Learning Hochreiter (1998)
Figure 7: Standard LSTM Structure Shi et al. (2015)
To illustrate the structure, the BiLSTM model is presented in the Figure 8.
Figure 8: BiLSTM Structure Gan et al. (2021)
Mathematics on CRFs
As an undirected graph model conforming to Markov random field, the conditional distribution $P(Y|X,\lambda)$ of target tag sequence y is calculated based on observation sequence X. Given a clinical text sequence $X=<x_{1},x_{2},...,x_{n}>$, CRFs model aims to make tags on X with a sequence $X=<y_{1},y_{2},...,y_{n}>$ to get this outcome $t_{i}\in\{B,I,E,O\}$, where B E I O represent the start, the end, the inner and the outside of a TIMEx corresponding to $x_{i}$. Given under the condition of random variable X, if the conditional probability distribution of random variable Y $P(Y|X,\lambda)$ that constitutes a conditional random field, which meets the Markov property
$P(Y_{i}|X,Y_{1},Y_{2},...,Y_{n})=P(Y_{i}|X,Y_{i-1},Y_{i+1})$ (1)
The conditional probability can be calculated by the following formula
$P(Y|X)=\frac{1}{Z(X)}e^{\sum_{i=1}^{I}\sum_{k=1}^{K}}\lambda_{k}f_{k}(i,Y_{i},Y_{i-1},X)$ (2)
where Z(X)is the normalisation factor
$Z(X)=\sum_{Y}e^{\sum_{i=1}^{I}\sum_{k=1}^{K}}\lambda_{k}f_{k}(i,Y_{i},Y_{i-1},X)$ (3)
Given the location of the input sequence X and i, where $f_{k}(i,Y_{i},Y_{i-1},X)$ represents the first k eigenvalue values of mark $Y_{i}$ and mark $Y_{i-1}$ where $\lambda_{k}$ represents the feature weights. From the formula, conditional random field model to the algorithm to calculate the sequence through the forward - after the characteristics of different position expectation and conditional probability, then using maximum likelihood estimation of quasi-newton method to find the solution of the model parameters, such as the use of dynamic programming in the Viterbi algorithm for decoding, and the sequence of the output data for testing as shown in the Figure 9. Additionally, there are other possible NLP systems that can be used to extract various features including (1) MedLEE, (2) KnowledgeMap, (3) a Dictionary-based Semantic Tagger (DST).
Figure 9: CRF Structure Lafferty et al. (2001)
Suppose our NER task uses a BiLSTM model with some word embeddings (illustrated in Figure 8). The BiLSTM provides the logits per class for each word, but each prediction for a single word is independent of the predictions for its neighbours. Each token is independently classified without the model knowing the predicted classes of its neighbours.
More on CNNs
Figure 10: Convoluted Neural Networks Topology Saha
To the best of our knowledge, CNN performs admirably on NLP-related tasks like semantic parsing, NER, and Relation Extraction(RE), where deep learning has been efficiently and widely used due to the prevailing interests of academics in this area. ConvNets’ architecture (shown in Figure 10) mimics the human brain’s neuronal connectivity pattern and takes design cues from the Visual Cortex’s structure. Only some areas of the visual field, referred to as the receptive field, are capable of eliciting responses from individual neurons when they are stimulated. A combination of these fields overlaps to encompass the entire visual field. CNN is able to receive an image as input, assign importance (learnable weights and biases) to distinct aspects/objects in the picture, and differentiate between them. Comparatively, ConvNet requires substantially less pre-processing than other classification techniques. In contrast to primitive approaches, ConvNets are capable of learning these filters/characteristics given sufficient training. This concept works just as well in words and characters as it does in practice.
More on BERTs
As discussed in Section H: More Background and Literatures, for downstream tasks modelled by BERT, especially NLP tasks that include sentence text, firstly, the sentences need to be encoded independently and pre-trained through BERT’s self-attention mechanism. It is important to note that in the text classification task, we will need to test the sentence and the label L together and form the text - tag on; this text classification task can be directly into the sentence to the task. Finally, fine-tune the BERT model to find the most suitable parameters for the corresponding task.
Figure 11: BERT Pre-trained Process Shi et al. (2020)
D: Time Expressions and NER Task
Time expressions, medical events, and time relations identifications were performed on 600 clinical notes, and pathological texts from cancer patients Bethard et al. (2016). Conditional Random Fields (CRFs) have been found to be an effective method to detect Time Expressions and events Lin et al. (2013), based on which, we take the CRFs-based method as the recogniser to extract the text spans of TIMEX (e.g. ‘This February’, ‘24th March, 2020’) and EVENT (e.g. ‘Hospitalization from…’, ‘chemotherapy is indicated from… ’). CRFs model is a conditional probability model proposed in 2001 Lafferty et al. (2001). It combines the characteristics of the hidden Markov model (HMM) Eddy (2004) and Maximum Entropy Markov model (MEMM) McCallum et al. (2000), avoids the problem of label bias through global normalisation, and achieves good results in the task of NER Han et al. (2013).
The entity tags used in the i2b2 challenge track 2009 and 2012 have some differences which are shown in Table 9 and Table 10.
Table 9: Tags Used in the i2b2 2009 Dataset.
Tag Meaning Example
m medication Percocet
do dosage 3.2mg
f frequency twice a day
mo mode (/route of administration) Mode: “nm” (not “Tablet”)
du duration 10-day course
r reasons Dizziness
Table 10: Tags Used in the i2b2 2012 Dataset.
Tag Meaning Example
CLINICAL_DEPT clinical department emergency room
EVIDENTIAL events that have an ‘evidential’ nature CT shows
TEST clinical tests CT
PROBLEM symptoms sickness
TREATMENT medications, surgeries and other procedures Levaquin
OCCURRENCE the default value for other event types He was readmitted for
E: Dataset and Metrics
Despite the popularity and importance of clinical free text mining research, given the sensitivity of patients Schriever (2014), there have not been many public data sets available, and significant challenges are present when researchers are extracting useful information from clinical free text automatically. The EHRs data for this work mostly comes from National NLP Clinical Challenges, “n2c2” formerly called “i2b2”, which are challenges and workshops designed to find ‘the potential of existing clinical records to yield insights that directly impact healthcare improvement’.
To deliver and evaluate the performance of the temporal information extraction system properly, the effectiveness metrics including accuracy, precision, recall, and F-score Han (2014) about the system will be analysed on the training data and test data. To evaluate the machine learning-based classifier, the confusion matrix is widely applied with True Positives indicating an outcome where the model correctly predicts the positive class (TP), True Negatives indicating an outcome where the model correctly predicts the negative class (TN), False Positives indicating an outcome where the model incorrectly predicts the positive class (FP), and False Negatives indicating a result where the model incorrectly predicts the negative category (FN) as measures Manning and Schutze (1999). To make acknowledge whether our system, when finished, would facilitate the work in hospitals, it can be sent to a medical agency to conduct a trial on actual situations and give feedback on it.
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
$Precision=\frac{TP}{TP+FP}$
$Recall=\frac{TP}{TP+FN}$
$F1=\frac{2*Precision*Recall}{Precision+Recall}$
$F1=\frac{2*TP}{2*TP+FP+FN}$
In terms of interpretation, extraction results will be demonstrated in a structured table or CSV shape file with a comprehensive explanation and figure presentation function if necessary. Considering that there are multiple categories of labels, two different kinds of evaluation criteria are applied: 1) macro average evaluation, the function to compute precision, recall and F1-score, for each label and returns the average without considering the proportion for each label in the dataset; and 2) weighted average evaluation, the function to compute precision, recall and F1-score for each label, and returns the average considering the proportion for each label in the dataset.
Data Statistics
The n2c2 dataset (both 2009 and 2012) is composed of clinical notes that have been de-identified. Their annotated formats are not consistent (Figure 12 and Figure 13).
Figure 12: i2b2 2009 Dataset Example
Figure 13: i2b2 2012 Dataset Example
It could be not very pleasant if the chosen model is expecting a different format of training data. A widely used training dataset format is CoNLL Nivre et al. (2007) for NLP NER tasks. In this data presentation, each token (in this case, an individual word or punctuation mark) sits on its own line. The entire tokenised document is presented in that first column. The entity tag is in the last column. A custom parser is required to transform the data from n2c2’s entity-only, offset-based annotation format into CoNLL’s all-token, table-based format. We applied this parser to convert the data from the n2c2 2009 dataset into CoNLL’s format, which data structure consists of sentence row number, separate words (tokens), POS tag and label (as shown in Figure 14). However, there could be slight differences among data from different tracks; this parser could be used on a different dataset.
Figure 14: CoNLL Format
To better understand the dataset, some statistical analysis was taken. As shown in Figure 15, with BIO tagging format, the i2b2 challenge 2009 track dataset labels consist of 13 categories: B-m, I-m, B-do, I-do, B-f, I-f, B-r, I-r, B-mo, I-mo, B-du, and I-du, among which the B-m tag is in the majority of this dataset. A re-sample was conducted to erase the imbalance of the data sample (cutting labels more than 3000 down to 3000).
Figure 15: 2009 Dataset Labels Statistical Analysis
Then, through the statistical analysis of the sentence length of the input text data, the expected value of the normal distribution of the sentence length is taken as the target length of the sentence input in the subsequent model. It can be seen in Figure 16 that the majority of the sentence length is 10-20 words.
Figure 16: 2009 Word Count Statistical Analysis
Similarly, the same work was conducted on the n2c2 challenge 2012 dataset. Most sentences are between 5 and 25 words in length with wider range distribution than the 2009 dataset, as seen in Figure 17.
Figure 17: 2012 Word Count Statistical Analysis
F: Model Analysis
BiLSTM-CRF
Each classification decision is conditionally independent when a simple BiLSTM is followed by a classifier. The linear-chain CRF explicitly describes dependencies between labels as a table, including transition scores between all possible label pairs.
If the labels adhere to a tight internal syntax, the CRF can acquire this information very quickly. There are numerous methods for encoding the output of NER. Still, they typically encode at least the Beginning, Inside, and Outside of an entity, which must be in a syntactically determined order. CRF model will recognise very soon that it is impossible for I-m to follow O, as it must always follow B-m.
As they are conditionally independent, BiLSTM may be uncertain as to whether it should place B-m on a drug or one dosage unit later and, therefore, output both. CRF layer that is aware that this is unlikely and enforces the tags’ intrinsic logic would output B-do, I-do.
LSTM is able to add or remove information to the state unit, which is controlled by structures called gates. Followed by a soft-max layer, although BiLSTM learns adequate context information in the model training stage, the output probabilities of soft-max are independent of each other, and there is no mutual dependence between the output values. Soft-max only selects the label corresponding to the current optimal solution as the output at each step, which cannot avoid the output of wrong labels. For example, arguably, I-m always occurs following B-m instead of B-r. However, with soft-max, this type of apparent mis-classification happens a lot. To solve this problem, following BiLSTM, CRF as the output layer, take account into the order information between labels, so as to obtain more accurate labels through probability calculation, shown in Figure 18.
Figure 18: The Model Structure of BiLSTM+CRF We Re-Implemented
CRF directly models the conditional probability distribution $P(Y|X)$ combining the advantages of graph models to predict multivariate outputs y with a large number of input features x. The linear chain CRF built on Bi-LSTM can effectively model several hard constraints containing the dependencies of output labels.
BiLSTM-CRF Evaluation
The network structure including parameter values of the model for the i2b2 challenge 2009 dataset NER task is shown in Figure 19. Several groups of control experiments are carried out under this model. The main parameters of the final model are shown in Table 11, and the optimal performance on the test set is obtained after 4 epoch training.
Figure 19: Network of BiLSTM Model on the 2009 Dataset
parameter value
DENSE_EMBEDDING 50
LSTM_UNITS 50
LSTM_DROPOUT 0.2
DENSE_UNITS 100
BATCH_SIZE 256
MAX_EPOCHS 10
LEARNING_RATE 0.001
Table 11: Model Parameters of BiLSTM+CRF for the 2009 Dataset
With 70% of the training dataset, 15% validation dataset and 15% test dataset setting, the training process for NER task on the i2b2 2009 dataset is shown in Figure 20 including Loss function and CRF accuracy learning curves.
Figure 20: Training Process on 2009 Dataset
The structure of the model for the i2b2 challenge 2012 dataset NER task is shown in Figure 21. Several groups of control experiments are carried out under this model as well. The main parameters of the final model are shown in Table 12, and the optimal performance on the test set is obtained after 10 epoch training.
Figure 21: Network of BiLSTM Model on the 2012 Dataset
parameter value
DENSE_EMBEDDING 50
LSTM_UNITS 50
LSTM_DROPOUT 0.2
DENSE_UNITS 100
BATCH_SIZE 256
MAX_EPOCHS 30
LEARNING_RATE 0.0001
Table 12: Model Parameters of BiLSTM+CRF for the 2012 Dataset
With 70% of training dataset, 15% validation dataset and 15% test dataset, the training process for NER task on the i2b2 2012 dataset is shown in Figure 22.
Figure 22: Training Process on the 2012 Dataset
CNN-BiLSTM
The seven layers in this model and their functions:
  1. 1.
    Character embedding layer, which maps the input of characters to 30 dimensional embedding.
  2. 2.
    Dropout layer (0.5), which mitigates the effect of overfitting
  3. 3.
    1D convolutional layer, which transforms the character dimension size into 1
  4. 4.
    Word embedding layer, which maps the words into 50 dimensional embedding vectors with Glove (glove.6B.50d)
  5. 5.
    Concatenation layer, which combines processed character-level, word-level and casing data into a vector of 80 dimensions
  6. 6.
    BiLSTM layer: using the merged word vector sequence as input, the spatial semantic modelling of the preceding and subsequent text information was carried out, and the bidirectional semantic dependence of word vectors was captured. The high-level feature expression of the context information of medical records was further constructed
  7. 7.
    Dense output layer, which applies softmax function for prediction.
The model parameters and structure of CNN-BiLSTM is displayed in Table 13 and Figure 23, with codes adapted from ‘github.com/kamalkraj’ [^2] https://github.com/kamalkraj/Named-Entity-Recognition-with-Bidirectional
-LSTM-CNNs
parameter value
DENSE_EMBEDDING 30
DROPOUT 0.5
WORDS_EMBEDDING 50
DROPOUT_RECURRENT 0.25
LSTM_STATE_SIZE 200
CONV_SIZE 3
LEARNING_RATE 0.0105
OPTIMIZER Nadam()
Table 13: Model Parameters of BiLSTM+CNN for the 2009 Dataset
Figure 23: Network of CNN-BiLSTM Model on the 2009 Dataset Chiu and Nichols (2016)
BERT-CNN
Code adapted from open source projects of UmlsBERT and MedicalRelationExtraction, we trained this BERT+CNN model based on the parameters on the i2b2 challenge 2012 dataset.
The parameters in BERT-CNN model are set as in Table 14 and the model structure is presented in Figure 24.
Figure 24: The Architecture of BERT-CNN Model We Re-Implemented
parameter value
BERT_MODEL $Bio-BERT_{base}$
DROPOUT 0.5
EPOCH 5
DROPOUT 0.1
MAX_SENTENCE_LENGTH 512
HIDDEN_SIZE 768
LEARNING_RATE 1e-05
Table 14: Model Parameters of BERT+CNN
Post-Processing
In order to apply the prediction results of the NER task for medication recognition and the RE task for temporal relation extraction between medication event and date on medication usage status, we design a set of rules that determine whether a medication is in use. We take the date of Admission and Discharge as a time-frame and take the relation between the time-frame and medication into consideration. The dataset texts come from de-identified discharge summaries (HOSPITAL COURSE), which give the treatment information of patients during hospital time. The temporal relation on medication is categorised into several situations for ‘in use’ and ‘not’ as below.
  1. 1.
    If the relation between Admission Date and a medication AFTER and the relation between Discharge Date and the medication is BEFORE or OVERLAP, it means the medication is in use.
  2. 2.
    If the relation between Admission Date and a medication OVERLAP and the relation between Discharge Date and the medication is OVERLAP, it means the medication is in use.
  3. 3.
    If the relation between a date except for Admission and Discharge Date which is between them and a medication is OVERLAP and the relation between Discharge Date and the medication is BEFORE or OVERLAP, it means the medication is in use.
  4. 4.
    Otherwise it is not.
With this set of rules, we manually checked the classification results, showing that nearly three out of ten can be classified correctly. For example, using the following text:
Admission Date : [R] 2015-09-14 [R] Discharge Date : [R] 2015-09-19 [R] Service : NEONATOLOGY HISTORY OF PRESENT ILLNESS : The patient is a 3285 gm infant born at 37 5/7 weeks to a 21 year old G3 P1 now 2 mother with prenatal screens as follows : O positive , antibody negative , hepatitis B surface antigen negative , RPR nonreactive , GBS negative . Unremarkable pregnancy except for minor fullness of the left renal pelvis reported during the week prior to delivery . Past OB history remarkable for postpartum depression . Mother was admitted in labor . Baby was delivered by repeat C-section with rupture of membranes at delivery . Apgars were 8 and 9 . CMED CSRU staff was called about 30 minutes of age for grunting , flaring and retractions and the baby was admitted to the CMED CSRU . HOSPITAL COURSE: 1. Respiratory . The Athol Memorial Hospital hospital course was initially consistent with transient tachypnea of the newborn . Chest x-ray revealed mild streakiness of the lung fields . He was initially placed on [L] nasal cannula [L]
The predicted relation between ’nasal cannula’ and Discharge Date is BEFORE, and the predicted relation between ’nasal cannula’ and Admission Date is AFTER, which indicates that the medication treatment is in use according to the rules. Table 15 shows some correct classifications.
id medication status
IN USE OUT OF USE
23 Enfamil
2015-09-14-
2015-09-19
2015-09-19 - UNKNOWN
23 Antibiotics
UNKNOWN-
2015-09-19
2015-09-19 - UNKNOWN
151 Diuretics 1993-03-26 - 1993-04-03 1993-04-03 - 1993-06-13
Table 15: Medication Status Classification Structured Output
G: Hardware and Computation
Considering the computation resources we possess, in order to meet with the computational ability of the models’ needs, the hardware and software used will be discussed in this section. Given the complex progress of this project, we have done the data pre-processing with our personal computer (configurations shown in Table 16).
Hardware Test Environment
Central Processing Unit (CPU) Intel Core i5 1.6GHz
Random Access Memory (RAM) 16GB 2133MHz DDR3
Graphics Processing Unit (GPU) Intel UHD Graphics 617 1536 MB
Disk Memory 256GB
Table 16: Hardware and Computational Testing Environment
The modelling in this project is a relatively large amount of computing resources-consuming process. It will take more than a week if we train the model on our personal computer, in which case, a Google virtual machine platform (Google Co-Lab) and Microsoft Azure Machine Learning Studio are used to train and test models. All the codes were run in Ubuntu 18.04, macOS Monterey 12.1, Python version 3.8, and Tensorflow version 2.1.4.
H: More Background and Literatures
Researchers have done significant studies on clinical text mining in the past few years. These research findings facilitate data-driven projects in hospitals (help doctors make better diagnoses and treat patients, and reduce the unnecessary process of drug use history inquiries). This section discusses the existing methods and research results we have surveyed in the extraction of temporal information. Typically, information extraction consists of Named Entity Recognition (NER) and Relation Extraction (RE). The research methods of NER and RE can be divided into rule-based and statistics-based methods (including traditional machine learning-based and deep learning-based), and the two methods can also be combined. In terms of accuracy, rule-based methods perform better than traditional statistics-based methods. However, with the development of deep learning neural networks, the performance of methods based on deep learning exceed that of rule-based methods Pustejovsky et al. (2003). In many cases, the establishment of rules is often related to the language environment and must be completed with the cooperation of experienced language experts. In contrast, statistics-based methods do not require extensive linguistic knowledge, which is acquired through machine learning and has good portability. The model only needs to be retrained during transplantation on a new corpus, but statistics-based methods require a large corpus for learning. So there are advantages and disadvantages to both approaches.
5.1 Rule Based Methods
Ruled-based methods, which require experts with specific specialties to make a set of reasonable rules, are subjective and vary from person to person with different opinions towards this domain. Technically, rule-based methods often constitute a set of regular expressions. They are encoded by experts with domain knowledge and experience based on linguistic analysis, which requires great time and effort. For NER rule-based methods, rules often based on the named entity itself and the context of a situation are generally formulated by linguists. Compared with other language scripts (such as Chinese), it is relatively easy for English corpora to name entities because English name entities usually start with a capital letter, such as personal name, organisation name, etc., all start with a capital letter. If the named entity is identified, then the corresponding rules can be used to determine which type the named entity belongs to. In various corpora, the rules are constantly improved through the error analysis of rule-based experimental results until a certain recognition accuracy is achieved. We list some relevant work below.
Strötgen and Gertz (2013) built a rule-based system pattern called HeidelTime that expects part-of-speech(POS) tagged sentences and user custom parameters determining what kinds of temporal expressions are specified to extract temporal expressions and obtained high-quality results, which are 90%, 82% and 86% in Precision, Recall, and F1-score respectively. Reeves et al. (2012) used handcrafted rules to annotate temporal expressions (TIMEx). They considered each TIMEx as a tuple $t_{i}\in\{date;time;duration\}$. Basically, they made use of regular expressions to construct patterns to match sentences with similar expressions using the designed rules and achieved at least an increase of 5% in the precision score.
In general, rule-based methods have been widely used in the last 2 decades and some of them Roberts et al. (2013); Lin et al. (2013) achieved remarkable results in Medical terminology Peek et al. (2013) and TIMEx Hao et al. (2018) extraction tasks. Language Wise: Hao et al. (2018) first produced a set of rules based on heuristic theory and specific temporal features training on clinical texts to generate structural patterns that work on other pattern-based temporal relations. This system achieved 0.916 of F1 score in extracting temporal expressions on Chinese clinical texts, which continually has influence over this research field. Rule-based methods can perform well in event recognition; however, we can not ignore that developing a comprehensive and heuristic set of rules is a rather time-consuming and difficult task. These rules often depend on specific languages, domains and text formats, which are difficult to achieve a broad language coverage. Therefore, it is difficult for humans, even experts, to consider the correct rules thoroughly in all aspects. Wang et al. (2012) who generated rules based on syntactic information, including dependency types and conjunctions for information extraction from biomedical literature, achieving an increase by 3%-5% than Coulet et al. (2010). In order to investigate whether the existing medical terminologies work on Sweden language clinical texts as well as it does on English language texts and different techniques for pre-processing that might affect the performance of the corresponding system, Skeppstedt et al. (2012) proposed a system using different pre-processing methods that evaluated existing systems (including SNOMED-CT Donnelly et al. (2006)) for Sweden clinical texts. The main rule of this system for selecting annotation categories is to annotate words into categories that perceive them in clinical reality as described in the text. Thus, the same word or expression may be a discovery in one case and a disease in another, and it is annotated according to context. Hypertension, for example, can be seen as a symptom of multiple diseases, but it is also a disease in its own right, with its own underlying pathological processes. With this system, they achieved a precision of 0.74 and a recall of 0.80 on the body structure entity. Uzuner et al. (2010) built a system to extract medical information (body parts, anatomical parts, surgery reason, disease) for automated clinical data processing. Uzuner et al. (2010) proposed two apps designed to select data from Polish-language medical documents: mammogram reports and electronic records of diabetics. For each domain (body part, disease or medication), they defined a set of appropriate rules that enabled them to identify its value. For example, the syntax for the mammogram report contains 190 rules, while the syntax for the diabetes field contains 150 rules. In the case of mammography, the rules cover almost the entire report, whereas, in the case of diabetes, they target only a subset of information. This system obtained significant results in precision and recall achieving more than 80% in most templates.
Another widely applied technical framework that facilitates users to build and develop natural language processing tasks (NER and IE) called GATE Cunningham H (2002), developed by the Natural Language Processing Group at the University of Sheffield with funding from the Engineering and Sciences Research Council. It is an open-source natural language processing platform from the UK. The three main purposes of GATE design are: 1. Provide a technical framework for software for various natural language processing tasks. 2. Provide component and class library services, which can be embedded in the natural language processing software. 3. Provide a development platform for a natural language processing system, allowing users to debug in the visual interface.
Despite all the hard work from other researchers, there still exists a vague definition regarding what could be a perfect rule as a system to extract name entities or relations between mentions. Besides some well-done research in the English language clinical free text information extraction, rule-based information extraction is still an intractable challenge. Besides, in general, a set of rules generated from a specific domain are hard to be extended to apply to other domain Ji et al. (2019); thus, rule-based methods are no longer the first option for temporal information extraction. With the development of computation resources and machining learning, it allows researchers to build more complex models but only costs an acceptable time to run the model that could learn sufficient features and custom configuration to make one domain corpus distinguish from other fields.
5.2 Machine Learning-based Methods
Compared with computer vision and speech recognition, deep learning is a latecomer to natural language processing. On the one hand, because the text does not have digital features, the way of modelling text features is not mature. On the other hand, because the traditional classical network algorithm can not get the temporal features of text, it can not achieve good results in the application of text processing. Through the efforts of many scholars for more than ten years, the application of deep learning to natural language processing has become increasingly mature. Bengio et al. (2000) systematically proposed Neural Network Language Model (NNLM) and conducted in-depth research. One of the core ideas presented by NNLM is word vectors. With the proposal of a word vector, the neural network becomes more and more mature for the modelling. Prior to the invention of word vectors, researchers mainly extracted text properties using statistical techniques like One-hot Chren (1998) or TF-IDF Wu et al. (2008). Although these two approaches can partially represent text, their drawbacks are extremely clear. First of all, One-HOT and TF-IDF are essentially just statistics of how frequently words occur in the text. Since each dimension in the vectors produced by this statistical method has distinct properties, they are unable to convey the text’s semantic content. Second, each one-hot and TF-IDF dimension corresponds to a word in the lexicon. It is straightforward to get into excessive dimensions for a little larger corpus. It is unavoidable to sacrifice certain words in order to lessen the effects of excessive dimensionality, which can result in the input statement losing some information.
With the rapid development of artificial intelligence and machine learning, machine learning-based methods have been widely used in NLP. Machine learning generally falls into two categories: supervised learning and unsupervised learning. Supervised learning, of which application is used wider, consists of three main parts:
  1. 1.
    The extraction of feature to represent the data;
  2. 2.
    A model to work as a classifier;
  3. 3.
    Process of tuning the parameters of the model.
Basically, the main part that differs between machine learning-based (Categorized into supervised and unsupervised learning) methods is how the algorithm works for feature engineering. The difference between supervised and unsupervised feature selection is that one works with labelled classes, and the other does not. However, it could be very time-consuming work for people to label the raw data if doing supervised learning classification. The method of named entity recognition based on statistical machine learning has been deeply studied and applied in view of the shortcomings of dictionary and rule-based methods. Among various statistical machine learning methods, the support vector machine (SVM) is widely used. It plays a vital role in solving pattern recognition problems such as small samples, linear inseparability and high dimension. The method based on the Hidden Markov model (HMM) is also one of the common methods of named entity recognition in the medical field. Zhou et al. (2005) combined one SVM classifier and two HMM classifiers and found that different classifiers had different results on the dataset and could complement each other, and the F1 value on the genetic dataset reached 0.83. However, HMM is based on the assumption of independence. Strict independence assumption can not truly describe the information contained in the data sequence and can not understand the semantic relationship between the remote contexts of text sentences. The appearance of conditional random fields (CRFs) solves the independence assumption of the hidden Markov model, considers the context information, and solves the semantic connection between the remote contexts of text statements. Compared with feature-based methods, deep learning-based methods help discover hidden features automatically and automatically learn the required representation from the original input in an end-to-end manner. The application of neural networks in NER was first proposed by Collobert et al. (2011), who deployed a model architecture based on temporal convolutional neural networks. When common prior knowledge is incorporated, the resulting system outperforms the baseline using only word-level representations. Wu et al. (2015) used convolutional layers to generate global features represented by several global hidden nodes. The local and global features were then fed into a standard affine network to extract named entities in clinical medical records that obtained a high F1-score of 92.8% in Chinese EMRs. Aguilar et al. (2019) proposed a multi-task approach for NER that uses CNN to capture positive spelling features and words at a character level. For the syntactic and contextual information at the word level, such as part of speech and word embedding, the model is implemented through the LSTM model architecture.
Jiang and Chen addressed that, with the CRFs algorithm, this system achieved at least 0.8475 in the F-measure Jiang et al. (2011). For TLINKs detection, machine learning-based shows a remarkable effect on this classification task Chang et al. (2013). Focusing on CRFs and support vector machine (SVM), Xu et al. (2013) proposed a system with three sub-systems, including clinical events extraction, temporal information extraction and temporal relation extraction. In terms of prediction tasks, applying temporal information as a feature to improve the performance of the model has been proved a feasible way Pang et al. (2021). Theoretically, RNN can process any long sequence, but with the continuous accumulation of time series, the gradient will decay exponentially, which makes it difficult for RNN to record long-distance historical information, and its performance is also restricted. In order to solve this problem, Hochreiter and Schmidhuber (1997) proposed the concept of the long short-term memory network(LSTM) and theoretically proved that this structure can well solve the problem of gradient disappearance and explosion.Dang et al. (2018) has addressed a bidirectional Long Short-Term Memory (an advanced deep learning neural network) which achieved 87.62% F1 score for the gene/protein NER and 93.14 and 84.68% for the chemical and disease NER, respectively. In the field of NLP, the most commonly used deep learning algorithm is based on the deep structure of the recurrent neural network. The traditional neural network can not process the continuous input of natural language with time-series characteristics. The improvement of RNN is that it adds a loop pointing to itself so that the network can make use of the sequence characteristics of the input, Wu et al. (2017) compared CNN and RNN with baseline CRF in the i2b2 2010 clinical concept extraction corpus. The results showed that the performance of RNN was better than the best clinical named entity recognition system based on SSVM, and the F1 value was 0.86.
Theoretically, RNN can process arbitrarily long sequences, but with the continuous accumulation of time series, the gradient will be exponentially attenuated, which makes it difficult for RNN to record historical information within a long distance, and its performance is restricted. However, the effect of the BiLSTM and CRFs deep learning method is still limited by the quality of the training set. Therefore, people began to pay attention to how to obtain a priori semantic knowledge from a large number of unlabeled texts to enhance the semantic representation and proposed a transformer-based pre-training model, such as the Bert model. Some researchersKim and Lee (2020); Lee et al. (2020); Jiang et al. (2019) improved accuracy, recall, and F1 value by combining the Bert word embedding model with the conventional bilstm CRF model and taking into account the polysemy of a word in conjunction with the context. Bert, as we can see, has excellent semantic analysis skills. However, there is no separator between Chinese words, so Bert can only cover up words, not words when using the Chinese corpus for pre-training. Thus, the semantic representation generated by the pre-training is only at the word level, and the word level semantic representation cannot be obtained; that is, the word information cannot be obtained during the pre-training process. Recently, BERT, as a pre-trained Deep Bidirectional Transformers model for human language understanding, has been widely used in NLP related tasks, and a BERT-BiLSTM-CRFs Ma et al. (2022) model was proposed in extracting temporal information from social media messages. As a matter of fact, the latest BERT adaptation Rasmy et al. (2021) has introduced this adapted model Med-BERT with fine-tuning pre-trained on a structured EHR dataset which showed tremendous improvement in disease prediction tasks. Yuan et al. (2022) has designed the BiLSTM-CRFs model and BiGRU-Dual Attention model which achieved a significant F-1 score of 87.87% and 88.05% in the Archaeological site text domain. Despite the tremendous success of the BiLSTM-CRFs model in recognition performance in NER tasks and its ability to preserve long-term information, still, there is an unavoidable error in long sentence extraction, also called tagging inconsistency in Luo et al. (2018). In order to cope with this problem, based on the standard BiLSTM-CRFs algorithm, Ji et al. (2019) proposed a system, introducing attention mechanism to BiLSTM-CRFs, and conducted this system on Chinese EMRs, which achieved an F1-score of 90.82% higher than other work on the test dataset.
Compared to traditional deep learning methods that require a large amount of corpus for training, transfer learning can transfer knowledge from the source domain to the target domain by relaxing the condition that training instances and test instances obey independent and identical distribution, which has a tremendous positive impact on many fields that are difficult to make progress due to insufficient training data. Instance-based deep transfer learning refers to adding some samples in the source domain to the target domain as a supplement to the training samples in the target domain through special weight adjustment measures and assigning appropriate weights to these selected instances. It is based on the assumption that although there are differences between the source domain and the target domain, some instances in the source domain can be utilised by the target domain with a specific weight. Ling et al. (2008) proposed a TrAdaBoost system that deletes samples in the source domain that are not similar to the target domain with AdaBoost-based technology. AskTrAdaBoost is a fast algorithm invented by Yao and Doretto (2010), which can speed up the retraining of new target domains. Different from TrAdaBoost, it is designed for classification problems.
However, to achieve efficient and outstanding performance, deep learning networks need a lot of reasonable feature engineering work, which is expert-required and time-consuming. With BiLSTM, words or tokens in sentences can be represented by required dimensional vectors (also called word embedding) to be taken as input fed to a Conditional Random Fields (CRFs) model(used for classification). Apart from these embedding vectors, which contain contextual information, Part-of-Speech(POS), Chunks and Dependencies(Dep) are widely used as features to improve model performance Luo et al. (2018). Nevertheless, the extraction of some temporal phrases (e.g. next week, for 2 weeks) is still a problematic and prerequisite task.
5.3 Hybrid Approaches
Hybrid approaches interweave rule-based and machine learning-based methods to extract temporal information from clinical free texts. Chang, Yung-Chun and Dai addressed that their system performed better by 0.97% than those of single-stage based method on test dataset Chang et al. (2013). Chang et al. (2013) also found that their prediction system’s performance on TLINK AFTER is poorer than the other two categories because of the imbalance of the dataset. Lin et al. (2013) proposed a system of six main steps : (1) data pre-processing, (2) temporal tagging by HeidelTime Strötgen and Gertz (2013), (3) TIMEX and EVENT named-entity extraction , (4) classification, (5) clinical temporal normalization strategies, and (6) post-processing.
The rapid development of temporal information extraction methods, with the work of Pustejovsky Pustejovsky et al. (2003), has prevailed among many researches since 2003. This specification language TimeML Pustejovsky et al. (2003) takes into account the ordering of dependency existing in narratives that allows some of the temporal expressions to be decided later after interpretation, which makes a great contribution to question answering tasks. Tang et al. (2013)proposed a rule-based and machine learning-based combined method, and their system achieved top rank in The 2012 i2b2 NLP challenge. Their system contains three parts: Event extraction, Temporal expression extraction and time linking( TLink ) extraction. Separating a CRFs based classifier combined with SVM into two-stages, this classifier allows for extracting medical tests, problems and treatments in the first stage and the other events are extracted in the second stage achieving 0.9374 of precision 0.9013 of F1-score.
Yang et al. (2015) addressed a hybrid method for adding temporal information as an additional feature to build a model that ‘enhances the visualization accuracy of lung motion’. Atluri et al. (2018) has made great summaries on Spatio-Temporal Data Mining problems and methods. However, rule constraints inevitably require a lot of linguistic knowledge, which is not common among computer science researchers.
Different from traditional hybrid methods, Zhang et al. (2018) proposed a hybrid technique that applies CNN combined with traditional RNN, which takes advantage of the ability of CNN and RNN to learn potential features from biomedical literature. Then these learnt features are used to extract biomedical relations between proteins. Generally, this hybrid method, which takes dependency path and dependency word sequence as input, has shown significant performance in extracting relation in LLL Pyysalo (2008) corpus dataset that boosts a single CNN or RNN model effectively with 0.852 and 0.961 in F1-score and Recall, respectively.
Benefiting from the acceleration of research and development of computation, it takes less time for researchers to run and debug models. It gives more time for them to focus on what features contribute the most information to the model as well as the structure of Neural Networks. Basically, machine learning-based methods have prevailed in NER and relation extraction tasks. In addition, Cocos et al. (2017) used the advancements in deep learning technology to prove the benefits of deep neural network technology by comparing it to the conventional CRF model, finding that the former requires less intervention from artificial features than the latter does, hence allowing for more accuracy and recall. Deep learning can automatically extract word features, eliminate subjectivity in feature selection, and enhance recognition accuracy. In order to consider the context information, solve the semantic connection between the remote contexts of text sentences, and solve the gradient disappearance and explosion problem of the RNN model in the process of entity recognition, this paper applied the deep learning model BiLSTM-CRF to improve the accuracy of named entity recognition by using the context information.
More on MedTem Initiatives/Objectives
The primary purpose of this work is to extract temporal information related to diagnoses, drugs, and symptoms from free clinical texts as shown in the example of Figure 1 and convert it to structured data as in Table 17. For the situation when corresponding temporal information is missing, we use the tag ‘Unknown’.
Methotrexate
IN USE OFF
start = May 2019
stop = February 2020
start = February 2020
stop = Unknown
Table 17: Medication Temporal Information Extraction Result Example
This work focuses on building a system which is used to extract temporal information from texts facilitating the work in hospitals (e.g. help doctors attain a better picture of patients’ medical history) and research in the clinical text mining domain (e.g. with incorporating extracted temporal information as a feature for any other NLP tasks). To build up this system with the ability to extract temporal information automatically, there were several specific steps we carried out:
  • Review the literature and design the annotation schema for adding temporal features with available tools.
  • Apply state-of-the-art ruled-based and machining learning-based methods to recognise EVENTs and TIMEx (time expressions) mentions in texts and explore the possibility of improving them.
  • Extract features related to a DRUG event and classify these events into appropriate categories.
  • Investigate whether the model works for MEDICATION events (it did).
  • Link relations to their corresponding entity.
Finally, with the models we built, a structured temporal feature is extracted for further research and medicinal purposes.